# Testing identifying assumptions in Tobit Models

### v2

@@ Abstract.
This paper develops sharp testable implications for Tobit and IV-Tobit models’ identifying assumptions: linear index specification, (joint) normality of latent errors, and treatment (instrument) exogeneity and relevance. The new sharp testable equalities can detect all possible observable violations of the identifying conditions. We propose a testing procedure for the model’s validity using existing inference methods for intersection bounds. Simulation results suggest adequate test size for large samples and reasonable power in detecting violations of the exogeneity assumption and error structure. We review and propose new alternative paths to partially identify the parameters of interest under less restrictive assumptions. We revisit a study of married women’s labor supply in Lee (1995) to demonstrate the practical implementation and usefulness of the test.
The present version is as of January 17, 2025, the first version is from December 2022. All errors are ours. Email address: otavio.bartalotti@monash.edu. We want to thank participants at seminars in the 2024 Jornadas Anuales de Economía del Banco Central del Uruguay and RMIT University, the 2024 IAAE and Southern Economic Association meetings.
@@ 1. Introduction
Since the seminal work of Tobin (1958), Tobit models have earned attention in economics, business and social sciences. [^1] According to Google Scholar Tobin’s original paper has more than 10000 citations. Tobin (1958) analyzed household expenditure on durable goods using a regression model that specifically incorporated that expenditure (the dependent variable) cannot be negative. This approach is related to a broader class of censored or truncated regression models, depending on whether observations outside a specified range are lost or censored. When applied researchers are interested in modelling limited dependent variables, potentially with mass accumulation points, the Tobit family of models provides structure to identify parameters of interest, such as the average treatment effect (ATE). Identification relies on three primary sources: (i) instrument exogeneity (or exogeneity of the variable of interest itself), (ii) normality of the model’s latent variables, and in the case of an instrumental variable approach to endogeneity, (iii) the relevance condition for the instrument. While researchers recognize the model’s restrictive nature, it remains a valuable tool in the empirical literature.
In this paper, we develop a test for the validity of the Tobit model’s structure and assumptions, providing three main contributions to the literature. The first is to provide sharp testable equalities that can detect all possible observable violations of the Tobit model. Second, we propose a test for the validity of the Tobit model’s identifying assumptions using the sharp equalities that characterize the model to check its falsifiability. Following recent literature, we convert the equalities into conditional moment inequalities and implement the test by existing inferential methods from Chernozhukov, Lee, and Rosen (2013). The Tobit family of models involves continuous outcomes with accumulation points and potentially continuous treatment, implying a large number of moment equalities to be tested. This creates significant challenges to test implementation. We propose a discretization of the space of the treatment and the outcome that balances the computational requirements and data availability for different parts of their joint support. This simplifies the implementation, making it easy to compute and providing an asymptotically valid testing procedure.
@@ 1.1. Previous Literature
There is a vast literature related to testing the validity of Tobit models and their assumptions, which we contribute to. Most of the preceding work focuses on testing a particular assumption or feature of the model while maintaining other assumptions and structures as valid. Nelson (1981) constructs a Hausman-type test for misspecification of the classic Tobit model (that is, normality, linear index, and homoskedasticity) where the maximum likelihood estimates are compared with method of moment estimates. Nelson’s test compares the sample proportion of non-censored observations with the hypothesized probability of being non-censored in the Tobit model. Bera, Jarque, and Lee (1984) state that the test is equivalent to the Lagrange multiplier (LM) test of the Tobit model against Cragg’s model (Lin and Schmidt, 1984) and propose an alternative LM test for the normality assumption against other distributions of the Pearson family of distributions while the remaining assumptions are maintained. Newey (1987b) considers both exogenous and endogenous explanatory variables cases using symmetrically censored least squares estimators to construct specification tests of normality and homoskedasticity assumptions via a Hausman-type specification test. Holden (2004) examines several statistics proposed to test the normality assumption in the Tobit (censored regression) model and reinterpreted them as a version of the LM (score) test for a common null hypothesis. Reynolds and Shonkwiler (1991) use an information matrix misspecification test to detect violations of the distributional assumptions of the Tobit model.
@@
Other tests include Drukker (2002), which operationalized conditional moment tests developed by Newey (1985) and Tauchen (1985) to the case of misspecification of the distribution of the classic Tobit model. With a similar intuition to our framework, their test writes down conditional moment restrictions, which should have zero conditional expected values under the null. Since the model was estimated by maximum likelihood, the assumed data-generating process specifies the moments of disturbances conditional on the covariates to be the ones of a normal distribution. Drukker (2002) use these moment-based methods based on the third and fourth moments of the normal distribution. While intuitively similar, our procedure detects all possible violations of the model, not only those evident from deviations in the third and fourth moments. Smith and Blundell (1986) propose a test of the treatment variable’s exogeneity in the IV-Tobit Model by a control function approach exploiting the joint normality of the latent variables. Most of these approaches focus on testing a particular subset of assumptions or a specific class of alternatives to consider.
@@
The developments proposed in this paper consider all observable violations of the general Tobit model structure and its assumptions, serving as a useful test for empirical researchers constructing models for censored/truncated data. This work contributes to the growing literature on the testability of the identifying assumptions in various econometric models. Those include tests for the validity of assumptions in instrumental variable models, local average treatment effects, regression discontinuity designs, bivariate probit, encouragement designs, among others. (Mourifié and Wan, 2017; Kitagawa, 2015; Huber and Mellace, 2015; Gunsilius, 2021; Arai et al., 2022; Acerenza, Bartalotti, and Kédagni, 2023; Bai and Tabord-Meehan, 2024). Our testing procedures are connected to Kédagni and Mourifié (2020), which provides tests for assumptions related to the instrumental variables in the model, while we also consider the implications of the parametric structure on the outcome and latent error structure, which are inherent to Tobit models.
@@
The two papers closest to ours are Acerenza, Bartalotti, and Kédagni (2023) and Goff, Kédagni, and Wu (2024). The first focuses on bivariate probit models, leading to a finite set of moment equalities as testable restrictions. As mentioned above, this study builds upon their work by tackling the case of the Tobit family of estimators, which provides additional technical and practical challenges given the continuous nature of outcome and treatment, including accumulation points. Goff, Kédagni, and Wu (2024) considers separable parametric instrumental variable models that induce a different set of moment equalities to be tested while our approach focuses on Tobit and IV-Tobit models, which can potentially be non-separable. [^2] While Goff, Kédagni, and Wu (2024) briefly mention the possibility of extending their results to non-separable models, they obtain different equalities from ours and do not discuss identification or sharpness. Implementation would also differ from our approach.
@@
The remainder of the paper is organized as follows. Section 2 presents the model and identifying assumptions. Section 3 derives the sharp testable implications. Section 4 discusses heuristically the identification of the model’s parameters. Section 5 outlines the testing procedure. Section 6 include simulation evidence about the test’s size while discussions about power are relegated to the Appendix. Section 7 discusses how to relax the assumptions in case of rejection. Section 8 provides an empirical illustration revisiting the study of married women’s labor supply from Lee (1995) to demonstrate the practical implementation and usefulness of the test. Finally, Section 9 concludes. Additional results and details are collected in Appendices A-F.
@@ 2. Models
Let $Y=\max(0,Y^{*})$ be the observed outcome taking values in $\mathcal{Y}\subset\mathbb{R}^{+}$ with $Y^{*}$ a latent continuous dependent variable taking values in $\mathcal{Y^{*}}\subset\mathbb{R}$. Both the treatment of interest, $D$, and the instrumental variable, $Z$, can be discrete or continuous and take values in $\mathcal{D}\subset\mathbb{R}$ and $\mathcal{Z}\subset\mathbb{R}$, respectively. We normalize the coefficients by the standard deviation of the convenient unobservable structure for the exposition. [^4] Identification for the normalized and non-normalized parameters is discussed in section 4.
@@ 2.1. Classic Tobit model
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ \\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\end{array}\right.$ (2.4)
@@
$U$ is distributed as $N(0,1)$.
Assumption 1 states that the treatment of interest is independent of the model’s unobservables. Assumption 2 imposes that the latent error has a standard normal distribution. [^5] The error could take $N(0,\sigma^{2})$ but after scale normalizations we can model it as $N(0,1)$.
Despite its simplicity, the classic Tobit model can accommodate continuous or discrete treatment, and the normality assumption for the latent error could be replaced by other distributional assumptions that preserve the identification of the parameters.
@@ 2.2. IV Tobit model
The IV Tobit considers the case in which the treatment of interest is endogenous, but an instrumental variable is available to identify its effect. Let the IV Tobit model be: [^6] Results for the IV-Tobit model are presented for continuous $D$ but hold for a binary treatment, $D=1\{\gamma_{0}+\gamma_{1}Z-V>0\}$. In that case, the parameters are identified only up to scale.
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\alpha_{0}+\alpha_{1}D+U\\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.8)
Alternatively, in its reduced form representation:
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\beta_{0}+\beta_{1}Z+W\\ D&=&\gamma_{0}+\gamma_{1}Z+V\end{array}\right.$ (2.12)
Where $U,V$ are the latent structural error terms. To that model structure, restrictions on the joint distribution of the latent variables and their relationship to the instrumental variable are added.

### v3

@@ Abstract.
We develop sharp, testable implications for the identifying assumptions of Tobit and IV-Tobit models: linear index, (joint) normality of errors, treatment (instrument) exogeneity, and relevance. The new sharp testable equalities can detect all possible observable violations of the identifying conditions. The proposed test procedure for the model’s validity uses existing inference methods for intersection bounds. Simulations suggest adequate test size and power in detecting exogeneity and error structure violations. We review and propose alternatives to partially identify the parameters of interest under less restrictive assumptions. We revisit a study of married women’s labor supply in lee1995semi to demonstrate the test’s practical implementation.
The present version is as of December 15, 2025, the first version is from December 2022. All errors are ours. Email address: otavio.bartalotti@monash.edu. We want to thank participants at seminars in the 2024 Jornadas Anuales de Economía del Banco Central del Uruguay and RMIT University, the 2024 IAAE and Southern Economic Association meetings.
@@ 1. Introduction
Since the seminal work of tobin_estimation_1958, Tobit models have earned attention in economics, business and social sciences. [^1] According to Google Scholar Tobin’s original paper has more than 10000 citations. tobin_estimation_1958 analyzed household expenditure on durable goods using a regression model that specifically incorporated that expenditure (the dependent variable) cannot be negative. This approach is related to a broader class of censored or truncated regression models, depending on whether observations outside a specified range are lost or censored. When applied researchers are interested in modelling limited dependent variables, potentially with mass accumulation points, the Tobit family of models provides structure to identify parameters of interest, such as the average treatment effect (ATE). Identification relies on three primary sources: (i) instrument exogeneity (or exogeneity of the variable of interest itself), (ii) normality of the model’s latent variables, and, in the case of an instrumental variable approach to endogeneity, (iii) the relevance condition for the instrument. While researchers recognize the model’s restrictive nature, it remains a valuable tool in the empirical literature.
In this paper, we develop a test for the validity of the Tobit model’s structure and assumptions, providing three main contributions to the literature. The first is to provide the set of sharp testable equalities that can detect all possible observable violations of the Tobit model. Second, we propose a test for the validity of the Tobit model’s identifying assumptions using the sharp equalities that characterize the model to check its falsifiability. Following recent literature, we convert the equalities into conditional moment inequalities and implement the test by existing inferential methods from chernozhukov_intersection_2013. The Tobit model family is used for continuous outcomes with accumulation points. In the case of household expenditure, the outcome may exhibit a zero accumulation due to censoring. When evaluating causality, researchers may be interested in a treatment variable, not necessarily binary, which would result in a large number of moment equalities to be tested. This creates significant challenges to test implementation. We propose a discretization of the space of the treatment and the outcome that balances the computational requirements and data availability for different parts of their joint support. This simplifies the implementation, making it easy to compute and providing an asymptotically valid testing procedure.
@@ 1.1. Previous Literature
There is a vast literature related to testing the validity of Tobit models and their assumptions, which we contribute to. Most of the preceding work focuses on testing a particular assumption or feature of the model while maintaining other assumptions and structures as valid. nelson_test_1981 constructs a Hausman-type test for misspecification of the classic Tobit model (that is, normality, linear index, and homoskedasticity) where the maximum likelihood estimates are compared with method of moment estimates. Nelson’s test compares the sample proportion of non-censored observations with the hypothesized probability of being non-censored in the Tobit model. bera_testing_1984 state that the test is equivalent to the Lagrange multiplier (LM) test of the Tobit model against Cragg’s model (lin_test_1984) and propose an alternative LM test for the normality assumption against other distributions of the Pearson family of distributions while the remaining assumptions are maintained. newey_specification_1987 considers both exogenous and endogenous explanatory variables cases using symmetrically censored least squares estimators to construct specification tests of normality and homoskedasticity assumptions via a Hausman-type specification test. holden_testing_2004 examines several statistics proposed to test the normality assumption in the Tobit (censored regression) model and reinterprets them as a version of the LM (score) test for a common null hypothesis. reynolds_testing_1991 use an information matrix misspecification test to detect violations of the distributional assumptions of the Tobit model.
@@
Other tests include drukker_bootstrapping_2002, which operationalized conditional moment tests developed by newey_maximum_1985 and tauchen_diagnostic_1985 to the case of misspecification of the distribution of the classic Tobit model. With a similar intuition to our framework, their test writes down conditional moment restrictions, which should have zero conditional expected values under the null. Since the model was estimated by maximum likelihood, the assumed data-generating process specifies the moments of disturbances conditional on the covariates to be the ones of a normal distribution. drukker_bootstrapping_2002 use these moment-based methods based on the third and fourth moments of the normal distribution. While intuitively similar, our procedure detects all possible violations of the model, not only those evident from deviations in the third and fourth moments. smith_exogeneity_1986 propose a test of the treatment variable’s exogeneity in the IV-Tobit Model by a control function approach exploiting the joint normality of the latent variables. Most of these approaches focus on testing a particular subset of assumptions or consider a specific class of alternatives.
@@
The developments proposed in this paper consider all observable violations of the general Tobit model structure and its assumptions, serving as a useful test for empirical researchers constructing models for censored/truncated data. This work contributes to the growing literature on the testability of the identifying assumptions in various econometric models. Those include sharp tests of the validity in instrumental variable models222gunsilius_nontestability_2021 proves that instrument validity cannot be tested in the case where the endogenous treatment is continuously distributed. and local average effects (mourifie_testing_2017; kitagawa_test_2015; huber_testing_2015; kedagni2020generalized), sharp tests of the validity of assumption in the regression discontinuity designs (arai_testing_2022), sharp tests of the assumptions of the bivariate probit (acerenza_testing_2023) and sharp tests in the context of encouragement designs (bai2024sharptestableimplicationsencouragement) among others. Our testing procedures are connected to kedagni2020generalized, which provides tests for assumptions related to the instrumental variables in the model, while we also consider the implications of the parametric structure on the outcome and latent error structure, which are inherent to Tobit models.
@@
The two papers closest to ours are acerenza_testing_2023 and goff2024testing. The first focuses on bivariate probit models, leading to a finite set of moment equalities as testable restrictions. As mentioned above, this study builds upon their work by tackling the case of the Tobit family of estimators, which provides additional technical and practical challenges given the continuous nature of outcome and treatment, including accumulation points. goff2024testing considers separable parametric instrumental variable models that induce a different set of sharp moment equalities to be tested while our approach focuses on Tobit and IV-Tobit models, which can potentially be non-separable. [^3] While goff2024testing briefly mentions the possibility of extending their results to non-separable models, they obtain different equalities from ours and do not discuss identification or sharpness in that context. Implementation would also differ from our approach.
@@
The remainder of the paper is organized as follows. Section 2 presents the model and identifying assumptions. Section 3 discusses heuristically the identification of the model’s parameters. Section 4 derives the sharp testable implications. Section 5 outlines the testing procedure. Section 6 include simulation evidence about the test’s size, while discussions about power are relegated to the Appendix. Section 7 discusses how to relax the assumptions in case of rejection. Section 8 provides an empirical illustration revisiting the study of married women’s labor supply from lee1995semi to demonstrate the practical implementation and usefulness of the test. Finally, Section 9 concludes. Additional results and details are collected in Appendices A-E.
@@ 2. Models
Let $Y=\max(0,Y^{*})$ be the observed outcome taking values in $\mathcal{Y}\subset\mathbb{R}^{+}$ with $Y^{*}$ a latent continuous dependent variable taking values in $\mathcal{Y^{*}}\subset\mathbb{R}$. Both the treatment of interest, $D$, and the instrumental variable, $Z$, can be discrete or continuous and take values in $\mathcal{D}\subset\mathbb{R}$ and $\mathcal{Z}\subset\mathbb{R}$, respectively. We consider the normalized coefficients by the standard deviation of the unobservable error term, which is convenient for the exposition. [^5] Identification for the normalized and non-normalized parameters is discussed in section 3.
@@ 2.1. Classic Tobit model
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ \\ Y^{*}&=&\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+U\end{array}\right$. (2.4)
@@
$U$ is distributed as $N(0,\sigma^{2})$.
Assumption 1 states that the treatment of interest is independent of the model’s unobservables. Assumption 2 imposes that the latent error has a normal distribution. To simplify some of the derivations later in the manuscript, we defined the scaled version the model for which the transformed error term would have distribution $N(0,1)$.
$\frac{Y^{*}}{\sigma}=\frac{\tilde{\alpha}_{0}}{\sigma}+\frac{\tilde{\alpha}_{1}}{{\sigma}}D+\frac{U}{\sigma}=\alpha_{0}+\alpha_{1}D+\frac{U}{\sigma}$
We refer to the standardized parameters without the tilde, thus for example $\alpha_{1}=\frac{\tilde{\alpha}_{1}}{\sigma}$.
Despite its simplicity, the classic Tobit model can accommodate continuous or discrete treatments, and the normality assumption for the latent error can be replaced with other distributional assumptions that preserve parameter identification.
@@ 2.2. IV Tobit model
The IV Tobit considers the case in which the treatment of interest is endogenous, but an instrumental variable is available to identify its effect. It can be traced back to heckman1977dummy; amemiya1979estimation; nelson1978specification among several approaches to consider censored variables with endogeneity. [^6] In Section 7, we discuss alternative models and how identification can still be achieved under different assumptions, including about the distribution of the latent error terms. Let the IV Tobit model be: [^7] Results for the IV-Tobit model are presented for continuous $D$ but hold for a binary treatment, $D=1\{\gamma_{0}+\gamma_{1}Z-V>0\}$. In that case, the parameters are identified only up to scale.
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\tilde{\alpha}_{0}+\tilde{\alpha}_{1}D+U\\ D&=&\tilde{\gamma}_{0}+\tilde{\gamma}_{1}Z+\tilde{V}\end{array}\right$. (2.8)
Where $U,\tilde{V}$ are the latent structural error terms. Alternatively, in its reduced form representation:
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\tilde{\beta}_{0}+\tilde{\beta}_{1}Z+\tilde{W}\\ D&=&\tilde{\gamma}_{0}+\tilde{\gamma}_{1}Z+\tilde{V}\end{array}\right$. (2.12)
Where $\tilde{\beta}_{0}=\tilde{\alpha}_{0}+\tilde{\alpha}_{1}\tilde{\gamma}_{0},\tilde{\beta}_{1}=\tilde{\alpha}_{1}\tilde{\gamma}_{1}$ and $\tilde{W},\tilde{V}$ are the reduce form error terms where $\tilde{W}=\tilde{\alpha}_{1}\tilde{V}+U$.
To that model structure, restrictions on the joint distribution of the latent variables and their relationship to the instrumental variable are added.
@@
Let $U,\tilde{V}$ follow a bivariate normal distribution with covariance $\rho_{U\tilde{V}}$, i.e., $\begin{pmatrix}U\\ \tilde{V}\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\ \end{pmatrix}$, and $\Sigma=\begin{pmatrix}\sigma^{2}_{U}&\rho_{U\tilde{V}}\\ \rho_{U\tilde{V}}&\sigma^{2}_{\tilde{V}}\end{pmatrix}$.